# Phantom matter: a challenging solution to the cosmological tensions

### v1

@@ 3 Phantom matter and the energy conditions of the cosmic fluids
Even though phantom matter was first proposed phenomenologically in the old composite cosmological model $\Lambda$XCDM [41, 42, 43], more recently it has been put forward on more formal grounds in [11], specifically within the context of the stringy RVM approach, see [10] for a review. Obviously, we need not provide technical details here and we refer the reader to the aforesaid references. However, we can at least describe the conceptual design of this picture. In a nutshell is the following: in a framework of a string-inspired cosmology with primordial gravitational waves and gravitational anomalies, which lead to dynamical inflation of RVM type without the need for ad hoc inflaton fields [19, 10], a crucial role is played by the fundamental axion field existing in the gravitational multiplet of string theory, viz. the Kalb-Ramond axion field $b(x)$, which couples to the gravitational Chern-Simons (gChS) term: $\sim b(x)\,R_{\mu\nu\rho\sigma}\,\widetilde{R}^{\mu\nu\rho\sigma}$ – here the tilde denoting the dual of the Riemann tensor. The KR axion, when combined with the gChS contribution, obey together a peculiar equation of state $p=-\rho$ with negative energy density $\rho<0$ (hence with positive pressure $p>0$), which defines the ‘phantom vacuum’[11]. This form of vacuum, however, is merely transitory until the gChS condensates $\langle R\,\tilde{R}\rangle$ become activated through the condensation of primordial gravitational waves (GW), thereby making possible a positively defined vacuum state with effective cosmological ‘constant’ $\Lambda(H)\sim\langle bR\tilde{R}\rangle$. At this juncture the total pressure and density involve the combined effect from all the contributions, what makes possible a normal vacuum state with $p_{\rm total}=-\rho_{\rm total}>0$, i.e. a standard de Sitter phase with positive density and negative pressure. The latter nevertheless may still be subdued to corrections owing to quantum effects. Interestingly enough, in such a framework the KR axion can be a candidate to Dark Matter, what provides an additional motivation for this cosmological picture since it could embrace the entire cosmic history.
@@
Thus, a phantom-matter dominated era can be present in the early cosmic evolution near a de Sitter phase, but we should emphasize that it could also reflourish in the late universe. As shown in  [52, 53, 54], owing to the generation of chiral matter, which will dominate the Universe at the exit from inflation, at large scales the universe can recover its FLRW background profile. This is because the chiral matter fields insure the cancellation of the gravitational anomalies during the FLRW regime, where they are indeed not observed as otherwise it would imply a glaring violation of general covariance. However, as soon as we are in the process of exiting the FLRW regime towards a new de Sitter era, we cannot exclude that in subsequent stages of the cosmic evolution the universe can be affected by the presence of lurking PM. The reason is that the chiral matter fields will get more and more diluted with the cosmic expansion and, then, owing to the incomplete cancellation of the gravitational anomalies, these may eventually re-surface near the current quasi-de Sitter era. It is therefore not inconceivable to think of the existence of phantom matter bubbles or domains tunneling now and then into our universe at relatively close redshift ranges before the ultimate de Sitter phase takes over in the universe. These bubbles of PM, being endowed with positive pressure $p>0$, could obviously foster a larger rate of the structure formation, a fact which might explain the overproduction of large scale structures at unexpected places and times deep in our past. As we shall next show, this ideology provides an excellent framework for a fit to the overall cosmological data which proves to be (far) better than that of the standard $\Lambda$CDM model. We refer once more the reader to [11] and the review [10] for the details underlying the PM scenario supporting our proposal for solving the tensions. Here we have limited ourselves to a summarized exposition and in the rest of this work we focus on exploring the phenomenological implications.
@@
Figure 1: Energy conditions for the cosmic fluids. The quintessence (Q) region ($-1<w<-1/3$ with $\rho>0$, $p<0$) is marked cross-hatched, and the conventional phantom region ($w\leq-1$ with $\rho>0$, $p<0$) is the blue sector indicated as P. The gray-dotted area corresponds to the peculiar “Phantom Matter” (PM) region: $w<-1$ with $\rho<0$ and $p>0$. Notice that PM satisfies the strong energy condition: $\rho+p\geq 0,\ \rho+3p\geq 0$ (all of the gray area) but not the weak energy condition: $\rho\geq 0,\ \rho+p\geq 0$ (shaded area, except the P and PM sectors). The DE component $X$ in our analysis behaves as PM, whereas the $Y$ component behaves as quintessence. The EoS line $w=-1$ marks off the classical vacuum. The RVM fulfills this EoS only approximately near our time owing to quantum effects [46]. See also [41], and [11] for the modern developments on PM.
Before closing this section and for the sake of a better contextualization of the PM option in the context of the class of energy conditions, we show in Fig. 1 some of the most common possibilities for the EoS satisfied by the cosmic fluids. In it, we highlight the phantom matter (PM) option. As can be seen, it is very different from the usual phantom DE, it lies actually in its antipodes. These two phantom-like components are therefore dramatically different.
@@ 4 Data and numerical analysis
In this work we constrain the various composite DE models described in section 2 making use of the following cosmological data sets:
  • The full Planck 2018 CMB temperature, polarization and lensing likelihoods [55].
@@
  • The SNIa contained in the Pantheon+ compilation [56], calibrated with the cosmic distance ladder measurements of the SH0ES Team [57]. We refer to this data set as SNIa+SH0ES for short.
@@
  • 33 data points on $H(z)$ in the redshift range $z\in[0.07,1.97]$ from cosmic chronometers (CCH) [58, 59, 60, 61, 62, 63, 64, 65, 66, 67], see Appendix A of [68]. We employ the corresponding full covariance matrix, as described in [69].
  • Transverse (aka angular or 2D) BAO data from Refs. [70, 71, 72, 73, 74]. This type of BAO data are claimed to be less subject to model-dependencies, since they are obtained without assuming any fiducial cosmology to convert angles and redshifts into distances to build the tracer map. These data points are extracted from the two-point angular correlation function or its Fourier transform. In anisotropic (or 3D) BAO analyses, instead, a fiducial cosmology is typically employed to construct the 3D map in redshift space, potentially introducing some model-dependence. Moreover, some studies have suggested that uncertainties may be underestimated by a factor of two in these analyses [75]. In contradistinction to the 3D BAO data, 2D BAO observations still give room for low-redshift solutions to the Hubble tension while respecting the constancy of the absolute magnitude of SNIa [39, 37].
@@
  • The data on large-scale structure at $z\lesssim 1.5$ from Refs. [76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86]. These are data points on the observable $f(z)\sigma_{8}(z)$, with $f(z)=-(1+z)d\ln\delta_{m}/dz$ the growth rate, and $\sigma_{8}(z)$ the rms mass fluctuations at a scale $R_{8}=8h^{-1}$ Mpc, see Table 3 of [49]. These data points, though, are taken by the observational groups using a fiducial cosmology with $h\sim 0.67$, which translates into measurements at a characteristic scale of $R\sim 12$ Mpc. Here, in contrast, we adhere to the reasoning and practice of Refs. [22, 23, 24], and we treat these observations as data points on $f(z)\sigma_{12}(z)$, using a Fourier-transformed top-hat window function $W(kR_{12})$ in the computation of $\sigma_{12}(z)$. The advantage, as emphasized by these authors, is that the scale $R_{12}=12$ Mpc is independent of the parameter $h$.
@@
We compute all the cosmological observables using a modified version of the Einstein-Boltzmann code CLASS [87, 88], and explore and constrain the parameter space of the various models using the Metropolis-Hastings algorithm [89, 90] implemented in MontePython [91, 92]. The resulting Monte Carlo Markov chains are analyzed with the Python code GetDist [93]. Our main numerical results are presented in Table 1 and in the triangle plots of Figs. 2-3. Supplementary information is provided in the Appendix, which will be commented in the main text..
@@
Table 1: Mean values and uncertainties at 68% CL obtained with the full data set CMB+CCH+SNIa+SH0ES+BAO+$f\sigma_{12}$. We show the best-fit values in brackets. We use the standard notations for the $\Lambda$CDM parameters. In the last three lines, we display the values of the minimum $\chi^{2}$, $\Delta$DIC and $\Delta$AIC, as defined in Eqs. (2) and (3), respectively. Positive values of $\Delta$DIC and $\Delta$AIC denote preference of the new models over the $\Lambda$CDM. As can be seen, the preference is very high.
In Table 1 we do not only list the mean of the various parameters, with their uncertainties and best-fit values, but also report the minimum values of $\chi^{2}$ obtained for each model and the difference between the deviance (DIC) [94] and Akaike (AIC) [95] information criteria found between the composite DE models and $\Lambda$CDM, which we treat as our fundamental benchmark model. These differences read $\Delta{\rm DIC}\equiv{\rm DIC}_{\Lambda{\rm CDM}}-{\rm DIC}_{i}$ and $\Delta{\rm AIC}\equiv{\rm AIC}_{\Lambda{\rm CDM}}-{\rm AIC}_{i}$, respectively, with $i$ referring to $w$XCDM or $\Lambda_{s}$. Both criteria penalize the use of additional parameters and can be regarded as a rigorous mathematical implementation of Occam’s razor. The DIC is defined as
@@
with $p_{D}=\overline{\chi^{2}}-\chi^{2}(\bar{\theta})$ the effective number of parameters in the model, $\overline{\chi^{2}}$ the mean value of the $\chi^{2}$ and $\bar{\theta}$ the mean of the parameters that are left free in the Monte Carlo analysis. Similarly, AIC is defined as
@@
with $n_{p}$ the number of parameters entering the fit. With our definition, a positive difference of these information criteria implies that the composite DE models perform better than the $\Lambda$CDM, whereas negative differences imply just the opposite. According to the usual jargon of the information criteria, if $20\leq\Delta\textrm{DIC}<$ 2 it is said that one finds weak evidence in favor of the new model under test (in this case, the composite DE models under scrutiny), as compared to the standard model. If $62\leq\Delta\textrm{DIC}<$ 6 we speak, instead, of positive evidence. If $106\leq\Delta\textrm{DIC}<$ 10, there is strong evidence in favor of the composite DE models, whilst if $\Delta\textrm{DIC}>10$ we can conclude that there is very strong evidence supporting the new model or models against the standard $\Lambda$CDM. An analogous consideration can be made using AIC, of course. DIC is considered to be a more accurate information criterion, since it incorporates the information encapsulated in the full Markov chains. However, for the sake of generality and to explicitly show the consistency between these two criteria, we display the results obtained from both criteria in our Table 1.
5 Discussion and conclusions
In this work, we have addressed a possible solution to the cosmological tensions within a simplified version of the old existing $\Lambda$XCDM framework [41, 42, 43] and in the light of new theoretical developments in the context of the stringy version of the running vacuum model (StRVM) [11, 10]. The entity $X$ that is involved in the $\Lambda$XCDM need not be a fundamental field but can display a phantom matter (PM) behaviour, which in the StRVM is predicted to appear in the transit to the de Sitter phase. The PM regime is only transitory and acts as a temporary stage with negative energy and positive pressure. It can be a pure anti-de Sitter (AdS) phase but in general its EoS $w_{X}$ is not necessarily $-1$, although it must satisfy $w_{X}\lesssim-1$. While in this Letter we have contented ourselves with a simplified version of the full $\Lambda$XCDM, which we have called the $w$XCDM, the analysis of the full model (in which the VED is running according to the QFT formulation) will be presented elsewhere[51]. Using provisionally the $w$XCDM model, we find that the PM is operative above a transition redshift $z_{t}\simeq 1.5$ (cf. Table 1). For $z>z_{t}$ the PM behavior is rapidly washed out since its energy density fraction becomes smaller and smaller in the past. And, on the other hand, its effect in the future ($z>z_{t}$) becomes erased by the progressive appearance of the Chern-Simons condensates, which finally establish a steady de Sitter regime[11]. For this reason, in practice the appearance of a PM phase is confined to a bubble of spacetime, which is left behind. However, in that space and time new structures can emerge, which are completely unsuspected in the context of the standard model. The PM bubbles which are met in the context of the stringy RVM approach originate as a result of a a “phantom vacuum” phase which emerges in the universe when we approach the de Sitter epoch. While the implications of this fact for the early universe were considered previously[11], here we have extended this phenomenon to the late universe. In both cases the PM bubbles are evanescent, in that they are not carried forward in time and only occur as tunneling attempts to establish the phantom vacuum[11] $p=-\rho>0$, until the Chern-Simons condensates of the StRVM neutralize these attempts and finally the late de Sitter era prevails. Below the redshift $z_{t}\simeq 1.5$ the transitory PM phase ‘disappears’ (it is left isolated in the past) in favor of the field $Y$ (the would-be running VED in the the full $\Lambda$XCDM approach) whose effective EoS is found to be of quintessence type ($w_{Y}\gtrsim-1$) in our analysis. This is the form of DE that lies in our most recent past and in fact reaches up to our time.

### v3

@@ 3 Phantom matter and the energy conditions of the cosmic fluids
Even though phantom matter was first proposed phenomenologically in the old composite cosmological model $\Lambda$XCDM (Grande et al., 2006, 2007, 2009), more recently it has been put forward on more formal grounds in (Mavromatos & Solà Peracaula, 2021b), specifically within the context of the stringy RVM approach, see (Mavromatos & Solà Peracaula, 2021a) for a review. Obviously, we need not provide technical details here and we refer the reader to the aforesaid references. However, we can at least describe the conceptual design of this picture. In a nutshell, it is the following: in a framework of a string-inspired cosmology with primordial gravitational waves and gravitational anomalies, which lead to dynamical inflation of RVM type without the need for ad hoc inflaton fields (Solà & Gómez-Valent, 2015; Mavromatos & Solà Peracaula, 2021a), a crucial role is played by the fundamental axion field existing in the gravitational multiplet of string theory, viz. the Kalb-Ramond (KR) axion field $b(x)$, which couples to the gravitational Chern-Simons (gChS) term: $\sim b(x)\,R_{\mu\nu\rho\sigma}\,\widetilde{R}^{\mu\nu\rho\sigma}\equiv b(x)R\,\tilde{R}$, where $\tilde{R}$ denotes the dual of the Riemann tensor. The KR axion, when combined with the gChS contribution, obey together a peculiar equation of state of vacuum type $p=-\rho$ but with negative energy density $\rho<0$ (hence with positive pressure $p>0$), which defines the ‘phantom vacuum’ (Mavromatos & Solà Peracaula, 2021b). This form of vacuum, however, is merely transitory until the gChS condensates $\langle R\,\tilde{R}\rangle$ become activated through the condensation of primordial gravitational waves (GW), thereby making possible a positive-definite vacuum state energy with effective cosmological ‘constant’ $\Lambda(H)\sim\langle bR\tilde{R}\rangle$. At this juncture, the total pressure and density already involve the combined effect from all the contributions, what makes possible a normal vacuum state with $\rho_{\rm total}>0$ and hence with $p_{\rm total}=-\rho_{\rm total}<0$, i.e. a standard de Sitter phase with positive energy density and negative pressure. The latter nevertheless may still be subdued to corrections owing to quantum effects. Interestingly enough, in such a framework the KR axion can be a candidate to Dark Matter, what provides an additional motivation for this cosmological picture since it could embrace the entire cosmic history.
@@
Thus, a phantom-matter dominated era can be present in the early cosmic evolution near a de Sitter phase, but we should emphasize that it could also reflourish in the late universe. As shown in  (Basilakos et al., 2019, 2020a, 2020b), owing to the generation of chiral matter, which will dominate the universe at the exit from inflation, at large scales the universe can recover its FLRW background profile. This is because the chiral matter fields insure the cancellation of the gravitational anomalies during the FLRW regime, where they are indeed not observed as otherwise it would imply a glaring violation of general covariance. However, as soon as we are in the process of exiting the FLRW regime towards a new de Sitter era, we cannot exclude that in successive stages of the cosmic evolution the universe can be affected by the presence of lurking PM. The reason is that the chiral matter fields will get more and more diluted with the cosmic expansion and, then, owing to the incomplete cancellation of the gravitational anomalies, these may eventually re-surface near the current quasi-de Sitter era. It is therefore not inconceivable to think of the existence of phantom matter bubbles or domains tunneling now and then into our universe at relatively close redshift ranges before the ultimate de Sitter phase takes over. These bubbles of PM, being endowed with positive pressure $p>0$, could obviously foster a larger rate of the structure formation, a fact which might explain the overproduction of large scale structures at unexpected places and times deep in our past. As we shall next show, this ideology provides an excellent framework for a fit to the overall cosmological data which proves to be (far) better than that of the standard $\Lambda$CDM model, provided that we exclusively focus on BAO 2D data in our fit. We refer once more the reader to (Mavromatos & Solà Peracaula, 2021b) and the review (Mavromatos & Solà Peracaula, 2021a) for the details underlying the PM scenario supporting our proposal for solving the tensions. See also (Mavromatos, 2022) and the recent developments (Dorlis et al., 2024; Mavromatos et al., 2024). Here we have limited ourselves to a very succinct exposition of the theoretical background and in the rest of this work we focus on exploring the phenomenological implications.
@@
Figure 1: EoS diagram for the energy conditions of the cosmic fluids. The quintessence (Q) region ($-1<w<-1/3$ with $\rho>0$, $p<0$) is marked cross-hatched, and the conventional phantom region ($w\leq-1$ with $\rho>0$, $p<0$) is the blue sector indicated as P. The gray-dotted area corresponds to the peculiar “Phantom Matter” (PM) region: $w<-1$ with $\rho<0$ and $p>0$. Notice that PM satisfies the strong energy condition: $\rho+p\geq 0,\ \rho+3p\geq 0$ (all of the gray area) but not the weak energy condition: $\rho\geq 0,\ \rho+p\geq 0$ (shaded area, except the P and PM sectors). The DE component $X$ in our analysis behaves as PM, whereas the $Y$ component behaves as quintessence. The EoS line $w=-1$ marks off the classical vacuum. The RVM fulfills this EoS only approximately near our time owing to quantum effects (Moreno-Pulido & Solà Peracaula, 2022). See also (Grande et al., 2006), and (Mavromatos & Solà Peracaula, 2021b) for the modern developments on PM.
Before closing this section and for the sake of a better contextualization of the PM option within the class of energy conditions, in Fig. 1 we show a few of the most common possibilities for the EoS of the cosmic fluids. In it, we particularly highlight the phantom DE (labelled P, in blue) and the phantom matter (PM) regions. The latter (marked gray-dotted) is far away from the usual phantom DE, it actually lies in its antipodes! These two phantom-like possibilities, both satisfying $w<-1$, are therefore dramatically different and must not be confused. It should also be emphasized that while phantom DE has been amply considered in the literature to describe different features of the DE, including a possible explanation for the $H_{0}$ and growth tensions in different frameworks (see the Introduction and references therein), in our approach phantom DE is not used at all. In point of fact, only PM is singled out in an optimal way within our specific proposal for a possible resolution of the cosmological tensions.
@@ 4 Data and numerical analysis
Let us now come back to the three composite DE models mentioned in Section 2. Leaving model i) for a separate study (Gómez-Valent & Solà Peracaula, 2024), in this work we constrain the DE models ii) and iii) making use of the following cosmological data sets:
  • The full Planck 2018 CMB temperature, polarization and lensing likelihoods (Aghanim et al., 2020).
@@
  • The SNIa contained in the Pantheon+ compilation (Brout et al., 2022), calibrated with the cosmic distance ladder measurements of the SH0ES Team (Riess et al., 2022). We will refer to this data set as SNIa+SH0ES for short.
@@
  • 33 data points on $H(z)$ in the redshift range $z\in[0.07,1.97]$ from cosmic chronometers (CCH) (Jiménez et al., 2003; Simon et al., 2005; Stern et al., 2010; Zhang et al., 2014; Moresco et al., 2012; Moresco, 2015; Moresco et al., 2016; Ratsimbazafy et al., 2017; Borghi et al., 2022; Tomasetti et al., 2023), see Appendix A of (Favale et al., 2024a). We employ the corresponding full covariance matrix, as described in (Moresco et al., 2020).
  • Transverse (aka angular or 2D) BAO data from Refs. (Carvalho et al., 2016; Alcaniz et al., 2017; Carvalho et al., 2020; de Carvalho et al., 2018, 2021). This type of BAO data are claimed to be less subject to model-dependencies, since they are obtained without assuming any fiducial cosmology to convert angles and redshifts into distances to build up the tracer map. The BAO 2D data points are extracted from the two-point angular correlation function or its Fourier transform. In anisotropic (or 3D) BAO analyses, instead, a fiducial cosmology (the standard $\Lambda$CDM model) is employed to construct the 3D map in redshift space, potentially introducing some model-dependency [^3] This is an important issue to keep in mind, which is in fact at the basis of our approach in the current presentation. It is known that the impact of the fiducial cosmology on the BAO measurements can be small (Carter et al., 2020; Heinesen et al., 2020; Bernal et al., 2020; Pan et al., 2024), but this is warranted only when such a cosmology is close to the true one. In fact, Anselmi et al. (2023) recently argued that these works do not explore a wide enough range of the fiducial parameter values allowed by the measured BAO distances, which are quite far away from the Planck best-fit $\Lambda$CDM model. In addition, it is claimed that by fitting the two-point correlation function while fixing at the same time the cosmological parameters (and also those entering the correlation function template) at fiducial values, might lead to an underestimation of the errors by roughly a factor of two, hence a rather significant potential effect (Anselmi et al., 2019).. It is also worth noticing that despite the fact of being constructed from the same parent catalogues of tracers, angular and anisotropic BAO data exhibit some degree of tension, see e.g. (Camarena & Marra, 2020) and, particularly, the very recent analysis of Favale et al. (2024b). In the last study, it is shown that upon excluding the radial component of BAO 3D (which is of course not present in BAO 2D) the ‘residual’ tensions left between the two BAO types can still attain in between $2\sigma$ to $4.6\sigma$, depending on the data set used. All in all, not surprisingly, these BAO data tensions may directly impinge in a significant way on the discussion of the cosmological tensions themselves, for the required solution to alleviate them may be conditioned to the specific BAO data type considered in the analysis (Gómez-Valent et al., 2024). We should also mention that, in contradistinction to the BAO 3D data, BAO 2D observations still offer room for low-redshift solutions to the Hubble tension while respecting the constancy of the absolute magnitude of SNIa (Akarsu et al., 2023; Gómez-Valent et al., 2024). Given, however, the persisting conflict between these two BAO types, our aim here is to remain as model-independent as possible in the light of the present knowledge, and for this reason we opt for using at this point the kind of BAO data which seem to be less subject to criticism at present (and in this sense that may be more trouble-free), to wit: the BAO 2D type. We leave the detailed comparison with the BAO 3D results for a more comprehensive exposition. A preview of such a comparative analysis will be nonetheless advanced in the conclusions (see Section 6) [^4] The extended discussion, including an explicit comparison between the results obtained from BAO 2D and BAO 3D data, as well as the computational details of the present work, will be disclosed in a forthcoming comprehensive study (Gómez-Valent & Solà Peracaula, 2024)..
@@
  • The data on large-scale structure (LSS) at $z\lesssim 1.5$ from Refs. (Guzzo et al., 2008; Song & Percival, 2009; Blake et al., 2011, 2013; Simpson et al., 2016; Gil-Marín et al., 2017; Hou et al., 2020; Said et al., 2020; Avila et al., 2021; Mohammad et al., 2018; Okumura et al., 2016). These are data points on the observable $f(z)\sigma_{8}(z)$, with $f(z)=-(1+z)d\ln\delta_{m}/dz$ the growth rate, $\delta_{m}=\delta\rho_{m}/\rho_{m}$ the matter density contrast, and $\sigma_{8}(z)$ the rms mass fluctuations at a scale $R_{8}=8h^{-1}$ Mpc, see Table 3 of (Solà Peracaula et al., 2023). These data points, though, are taken by the observational groups using a fiducial cosmology with $h\sim 0.67$, which translates into measurements at a characteristic scale of $R\sim 12$ Mpc. Here, in contrast, we adhere to the reasoning and practice of Refs. (Sánchez, 2020; Semenaite et al., 2022, 2023), and we treat these observations as data points on $f(z)\sigma_{12}(z)$, using a Fourier-transformed top-hat window function $W(kR_{12})$ in the computation of $\sigma_{12}(z)$. The advantage, as emphasized by these authors, is that the scale $R_{12}=12$ Mpc is independent of the parameter $h$.
@@
We compute all the cosmological observables using a modified version of the Einstein-Boltzmann code CLASS (Lesgourgues, 2011; Blas et al., 2011), and explore and constrain the parameter space of the various models using the Metropolis-Hastings algorithm (Metropolis et al., 1953; Hastings, 1970) implemented in MontePython (Audren et al., 2013; Brinckmann & Lesgourgues, 2019). The resulting Monte Carlo Markov chains are analyzed with the Python code GetDist (Lewis, 2019). Our main numerical results are presented in Table 1 and in the triangle plots of Figs. 2-3. Supplementary information on plots and tables is provided in the Appendix, which will be commented on in the main text.
@@
Table 1: Mean values and uncertainties at 68% CL obtained with the full data set CMB+CCH+SNIa+SH0ES+BAO+$f\sigma_{12}$. We show the best-fit values in brackets. We use the standard notations for the $\Lambda$CDM parameters. In the last three lines, we display the values of the minimum $\chi^{2}$, $\Delta$DIC and $\Delta$AIC, as defined in Eqs. (2) and (3), respectively. Positive values of $\Delta$DIC and $\Delta$AIC denote preference of the new models over the $\Lambda$CDM. As can be seen, the preference is extraordinarily high.
In Table 1 we do not only list the mean of the various parameters, with their uncertainties and best-fit values, but also report the minimum values of $\chi^{2}$ obtained for each model and the difference between the deviance (DIC) (Spiegelhalter et al., 2002) and Akaike (AIC) (Akaike, 1974) information criteria found between the composite DE models and $\Lambda$CDM, which we treat as our fundamental benchmark model. These differences read $\Delta{\rm DIC}\equiv{\rm DIC}_{\Lambda{\rm CDM}}-{\rm DIC}_{i}$ and $\Delta{\rm AIC}\equiv{\rm AIC}_{\Lambda{\rm CDM}}-{\rm AIC}_{i}$, respectively, with $i$ referring to $w$XCDM or $\Lambda_{s}$. Both criteria penalize the use of additional parameters and can be regarded as a rigorous mathematical implementation of Occam’s razor. The DIC is defined as
@@
with $p_{D}=\overline{\chi^{2}}-\chi^{2}(\bar{\theta})$ the effective number of parameters in the model, $\overline{\chi^{2}}$ the mean value of $\chi^{2}$, and $\bar{\theta}$ the mean of the parameters that are left free in the Monte Carlo analysis. Similarly, AIC is defined as
@@
with $n_{p}$ the number of free parameters entering the fit. We note that the above formula is a good approximation when the number of data points entering the fit is much larger than $n_{p}$, which is certainly the case here. With our definition, a positive difference of these information criteria implies that the composite DE models perform better than the $\Lambda$CDM, whereas negative differences imply just the opposite. According to the usual jargon of the information criteria, if $20\leq\Delta\textrm{DIC}<$ 2 it is said that one finds weak evidence in favor of the new model under test (in this case, the composite DE models under scrutiny), as compared to the standard model. If $62\leq\Delta\textrm{DIC}<$ 6 we speak, instead, of positive evidence. If $106\leq\Delta\textrm{DIC}<$ 10, there is strong evidence in favor of the composite DE models, whilst for $\Delta\textrm{DIC}>10$ we may legitimately conclude that there is very strong statistical evidence supporting the new model or models against the standard $\Lambda$CDM. Analogous considerations can be made using AIC, of course. DIC is considered to be a more accurate information criterion, since it incorporates the information encapsulated in the full Markov chains. However, for the sake of generality and to explicitly show the consistency between these two criteria, we display the results obtained from both criteria at the same time in our Table 1.
5 Discussion